# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ 6. Proofs
Let $F_{\varepsilon}({\boldsymbol{x}})$ be the polynomial matrix in Lemma 6.2 where $\varepsilon>0$ is defined in (17). Then, applying Theorem 6.1 to $F_{\varepsilon}({\boldsymbol{x}})$ yields $N_{1}\in\mathbb{N}$ and polynomial matrices $S^{\prime}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S^{\prime}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy that $S^{\prime}_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$, $S^{\prime}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$’s are SOS matrices, such that
@@
$L(\Gamma)\coloneqq\max_{{\boldsymbol{y}}\in\mathcal{Y}}\max_{{\boldsymbol{\alpha}}\in\mbox{supp}_{{\boldsymbol{x}}}(\Gamma)}\frac{{\boldsymbol{\alpha}}!}{\deg_{{\boldsymbol{x}}}(\Gamma)!}\|\Gamma_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\|.$
@@
$N_{2}\coloneqq\max\left\{\frac{\deg(\Gamma_{0})(\deg(\Gamma_{0})-1)L(\Gamma_{0})}{2\varepsilon}-\deg(\Gamma_{0}),\ \frac{\deg_{{\boldsymbol{x}}}(\Gamma)(\deg_{{\boldsymbol{x}}}(\Gamma)-1)L(\Gamma)}{2\varepsilon}-\deg_{{\boldsymbol{x}}}(\Gamma)\right\}.$
Let $S_{0}({\boldsymbol{x}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma_{0}$ and $S({\boldsymbol{x}},{\boldsymbol{y}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma$. Write
$S_{0}({\boldsymbol{x}})=\sum_{|{\boldsymbol{\alpha}}|=\deg S_{0}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\quad\text{ and }\quad S({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}.$

### v4

@@ 6. Proofs
Let $F_{\varepsilon}({\boldsymbol{x}})$ be the polynomial matrix in Lemma 6.2 where $\varepsilon>0$ is defined in (24). Then, applying Theorem 6.1 to $F_{\varepsilon}({\boldsymbol{x}})$ yields $N_{1}\in\mathbb{N}$ and polynomial matrices $^{\prime}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$^{\prime}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} and satisfy that $S^{\prime}_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$, $S^{\prime}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$’s are SOS matrices, such that
@@
$L(\Gamma)\coloneqq\max_{{\boldsymbol{y}}\in\mathcal{Y}}\max_{{\boldsymbol{\alpha}}\in\mbox{supp}_{{\boldsymbol{x}}}(\Gamma)}\frac{{\boldsymbol{\alpha}}!}{\deg_{{\boldsymbol{x}}}\Gamma!}\|\Gamma_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\|.$
@@
$N_{2}\coloneqq\max\left\{\frac{\deg\Gamma_{0}(\deg\Gamma_{0}-1)L(\Gamma_{0})}{2\varepsilon}-\deg\Gamma_{0},\ \frac{\deg_{{\boldsymbol{x}}}\Gamma(\deg_{{\boldsymbol{x}}}\Gamma-1)L(\Gamma)}{2\varepsilon}-\deg_{{\boldsymbol{x}}}\Gamma\right\}.$
Let $P_{0}({\boldsymbol{x}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma_{0}$ and $P({\boldsymbol{x}},{\boldsymbol{y}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma$. Write
$P_{0}({\boldsymbol{x}})=\sum_{|{\boldsymbol{\alpha}}|=\deg P_{0}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\quad\text{ and }\quad P({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}P}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}.$